OpenStudy (anonymous):

Solve the equation for x 1+ square root of x-2 = square root of x+3 The book says x=6 but I can't figure out how they got it. I keep going through the steps and the x's cancel out each other

OpenStudy (anonymous):



OpenStudy (anonymous):

I tried squaring both sides like my teacher told me, but the x's still cancel

OpenStudy (anonymous):

Sorry just got really confused. You're helping a ton, I'm just learning this stuff slow. First time I've ever seen anything like this!

OpenStudy (zzr0ck3r):

\(1+\sqrt{x-2}=\sqrt{x+3}\) square both sides \(1+2\sqrt{x-2}+x-2 = x+3\) so \(\sqrt{x-2}=\frac{x+3-x-1+2}{2}=2\) square both sides \(x-2=4\\x=6\)

OpenStudy (anonymous):

How come squaring it on the right produced all that stuff and on the left, it just came down with the square root symbol taken off?

OpenStudy (zzr0ck3r):

\(1+\sqrt{x-2}=\sqrt{x+3}\\so\\(1+\sqrt{x-2})^2=(\sqrt{x+3})^2\) we will deal with the left side first, foil on the left \(1^2+\sqrt{x-2}+\sqrt{x-2}+(\sqrt{x-2})^2=(\sqrt{x+3})^2\\=1+2\sqrt{x-2}+x-2=(\sqrt{x+3})^2\) now we deal with the right hand side \(1+2\sqrt{x-2}+x-2=x+3\) now collect everything not in the square root term to the right hand side \[2\sqrt{x-2}=x+3+2-x-1 = 4\] so we have \(2\sqrt{x-2}=4\) divide by the 2 \(\sqrt{x-2}=2\) now square both sides again \((\sqrt{x-2})^2=2^2\) this gives \(x-2=4\) so, finally..... \(x=6\)

OpenStudy (zzr0ck3r):

\((\sqrt{a})^2=a\) \(\)

OpenStudy (anonymous):

Thank you so much, I'm starting to finally understand it! I have a test on this tomorrow (most of them aren't anywhere as hard as this one) but now at least I get what to do for the steps. Great help thanks again